OpenStudy (anonymous):

Given f(x) = x+7/5 solve for f^-1(3). -20 2 8 10

OpenStudy (anonymous):

\[\huge f(x) = x+\frac{7}{5}\]is this your question?

OpenStudy (anonymous):

no it is x+7 ----- 5

OpenStudy (anonymous):

13

OpenStudy (anonymous):

its f of 3

OpenStudy (zale101):

f(x)=x+7/5 let f(x) be y y=x+7/5 Switch the x and y. x=y+7/5 5x=y+7 subtract 7 to both sides 5x-7=y f(3)=5(3)-7 ? lol sorry

OpenStudy (anonymous):

8 !

OpenStudy (anonymous):

Thank you for your help!

OpenStudy (zale101):

yes

OpenStudy (zale101):

np :)